The following three substances are placed together in thermal contact:

1) Mg = 1.0 kg gold at temperature Tg = −13 C

2) Ms = 2.0 kg silver at temperature Ts = +7.0 C

3) Mi = 0.01 kg ice at temperature Ti = 0.0 C

What is the final equilibrium temperature T ?

Data:

specific heat of ice: Ci = 2100

specific latent heat of fusion of ice/water: Lf = 3.35 × 10^5

specific heat of gold: Cg = 126

specific heat of silver: Cs = 234
Answer #1

Heat required for gold to change its temperature from т, =-13 С to  T=0degree C  is  H_g=M_gC_g(T-T_g)=1.0*126*(0+13)=1638,J

Heat given out by silver to change its temperature from т , -7 ° C to  T=0degree C  is  H, = M,C,(T,-Т) = 2.0 * 234 * (7-0) = 3276 J

Heat required for ice at T, = 0°C to change in to water at T=0degree C is

H_i=m_iL=0.01*3.35*10^5=3350,J

Out of the heat H3276J  given by silver, a part of it 1638 J is used to raise the temperature of gold from 13 ° C to 0°C and remaining heat 3276 - 1638 1638 J is used by a part of ice to change its state into water at 0°C.

A part of ice changes to water and the remaining will be left as ice 0°C.

Hence the equilibrium temperature of the system will be 0°C.